[Hallc_running] [New Logentry] Beam update

hszumila at jlab.org hszumila at jlab.org
Sun Feb 11 10:30:04 EST 2018

Logentry Text:
After the accelerator re-tuning seemed to be going very well yesterday morning, Dave Meekins came in and cooled down the target (both Loop 2 and Loop 3). Just after the start of swing shift, we received a call to expect beam soon in Hall C. Towards the end of swing, we started to receive beam. They got up to about 50-55 uA when testing the current, but A was not taking beam during that time. It was observed that the A and C beams were still not collinear. They started to work the problem back at the injector around midnight when they had the capture and quarter cryo modules trip off. As of this morning, that trip has been resolved and they are back looking at the injector. It will be a few hours before they start sending beam back into Hall C.

No run meeting today as the status is unchanged since Friday. We are back to full staffing of shifts, and I thank everyone for quickly adjusting to the changes yesterday. We will meet tomorrow, Monday. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3530400
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180211/ba7d8406/attachment.html>

More information about the Hallc_running mailing list